KyledaGreat:

The total revenue for Jane's Vacation Rentals is given as the function R(x)=100x−0.25x^2, where x is the number of villas rented. What number of villas rented produces the maximum revenue?

KyledaGreat:

@vocaloid

Vocaloid:

graphing would be the easiest way to solve this, simply graph and look where the vertex is alternate method: when you have a quadratic equation ax^2 + bx + c a is the coefficient of the x^2 term, b is the coefficient of the x term, c is the constant the x-coordinate of the vertex is x = -b/(2a)

KyledaGreat:

1 attachment
Vocaloid:

if you're using mathway: enter the function 100x−0.25x^2, find the maximum/minimum value, take the x-coordinate since that's the # of villas

KyledaGreat:

(200,10000)

Vocaloid:

good, and since it's asking for the # of villas, you would just take the x-coordinate 200

KyledaGreat:

The total cost of producing a type of truck is given by C(x)=17000−90x+0.2x^2, where x is the number of trucks produced. How many trucks should be produced to incur minimum cost?

Vocaloid:

similar logic, enter the function, look for the minimum, take the x-value

KyledaGreat:

(225,−8425)

Vocaloid:

good, and since x = number of trucks, you only want the x-coordinate 225

KyledaGreat:

Consider the following function. t(x)=(x−4)^2−9 Find the x-intercepts, if any. Express the intercept(s) as ordered pair(s). How many x-intercepts does this function have ? A. None B. One C. Two

KyledaGreat:

it looks like 2

Vocaloid:

good

KyledaGreat:

Consider the following function. t(x)=(x−4)^2−9 Find two points on the graph of the parabola other than the vertex and x-intercepts. A : ( , ) B : ( , )

Vocaloid:

it's in vertex form (x-h)^2 - k, where (h,k) is the vertex, so the vertex is (4,9) to find the x-intercepts, we set t(x) = 0 and solve for x 0 = (x-4)^2 - 9 9 = (x-4)^2 square root of both sides x-4 = 3 x - 4 = -3 so x = 7 or x = 1 ****however*** it is asking for points **other than** the vertex and x-intercepts, so you have to plug in any x-value ***Except*** 1, 4, and 7 so you could plug in x = 2 and x = 3 to get your two points

KyledaGreat:

x = 2 and x = 9 ?

Vocaloid:

yes, you could use x = 2 and x = 9, but you have to plug them into the function t(x)=(x−4)^2−9 to get the two points

Vocaloid:

calculate the value of the function at x = 2 and x = 9

KyledaGreat:

t(2)=(9−4)^2−9

Vocaloid:

good, keep going

KyledaGreat:

16

Vocaloid:

good, so (9,16) is one possible point repeat with x = 2

KyledaGreat:

okay, where i do i apply x = 2 in it ?

KyledaGreat:

and is (9, 16) for A ?

Vocaloid:

just like you plugged in x = 9, plug in x = 2 into the function. yes, you plug in point A as (9,16)

KyledaGreat:

oh i see what you mean

KyledaGreat:

23

Vocaloid:

check your arithmetic again t(x)=(x−4)^2−9 x = 2, so t(x) = (2-4)^2 - 9 = 4 - 9 = -5 so (2,-5) is the second point B

KyledaGreat:

Consider the following function. t(x)=(x−4)^2−9 Graph the parabola.

KyledaGreat:

1 attachment
Vocaloid:

going off the points we've calculated so far: the intercepts (1,0) and (7,0) and the two points we just calculated: (9,16) and (2,-5), plot these points, draw the parabola through them

KyledaGreat:

1 attachment
KyledaGreat:

i'm not sure how this can be plotted on the graph on mine

KyledaGreat:

Among all rectangles that have a perimeter of 84, find the dimensions of the one whose area is largest. Write your answers as fractions reduced to lowest terms.